Mrs. Cruz bought 30m of cloth consisting of linen and cotton. The linen per meter was P75 while the cost of cotton per meter was P25. If the total cost of the material was P1750, how many meters of each cloth did she buy?
Answer by CubeyThePenguin(3113)

L = meters of linen
C = meters of cloth

L + C = 30
75L + 25C = 1750

Multiply the first equation by 75.

75L + 75C = 2250
75L + 25C = 1750

Subtract to get 50C = 500, so C = 10 and L = 20.

Mrs. Cruz bought 10 meters of cotton and 20 meters of linen.